Solve the equation by using square root property.
(x+6)^2=-18

Answer by edjones(7569)

(x+6)^2=-18
x+6=+-sqrt(-18) Take square root of each side.
...=+-sqrt(-9*2)
...=+-3sqrt(2)i
x=-6+3sqrt(2)i, x=-6-3sqrt(2)i
.
Ed